a ball is thrown up in the air with a speed of 30 m/s. It returns to where it was thrown from after a time period of 6s. what is the maximum acceleration reached by this ball? what is the acceleration of this ball at this height?
Answer #1

V = 30m/s , t = 6s

Maximum height = H = V^2/2g = 30^2/2×9.8 = 900/19.6 = 45.92m

=) H = 45.92m

Here acceleration will be same through out the journey that is equal to 9.8m/s^2